[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Mar 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Mar 30 11:06:01 UTC 2017


HSU Shift Log for Thu Mar 30 2017 07:05:54 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Ridges and Tilefish working fairly well now.

AWIPS-2/NCP issues: GFS data is still delayed, with the 6Z arriving roughly  
an hour later than it should be.

Other Techincal issues: ATT still working the trouble ticket as of  
yesterday afternoon. All of the phones are unplugged so to hear it you  
would have to plug one into the power/phone lines. Now appears to be  
resolved.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170330/53422ce8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list